Find the ecentricity and equation of the asympotes of the hyperbola 4x^2-9y^2=9
1
Expert's answer
2020-07-28T18:58:56-0400

Given ,

"4x^2-9y^2=9\\\implies \frac{x^2}{(\frac{3}{2})^2}-\frac{y^2}{1^2}=1"

Thus, "a=\frac{3}{2},b=1" and since, eccentricity is

"e=\sqrt{1+(\frac{b}{a})^2}\\
\implies e=\frac{\sqrt{13}}{3}"

Asymptotes of hyperbola is given by

"\frac{x^2}{(\frac{3}{2})^2}-\frac{y^2}{1^2}=0\implies y=\pm\frac{2}{3}x"